# Do Code LLMs Do Static Analysis?

### v1

@@ 3.3 Static Analysis Tasks
In this section, we discuss three static analysis tasks in our experiments i.e. AST, data flow graph, and callgraph generation. We use different tools to generate the gold for each task. We use C/C++ and Java in our experiments.
@@
Abstract Syntax Tree Generation An Abstract Syntax Tree (AST) is a tree structure that captures the syntactic information of the source code. In addition to being a key intermediate representation in code compilation, ASTs are used to localize software bugs (Neamtiu et al., 2005), comprehend source code (Sun et al., 2023a; Welty, 1997), and in software evolution (Neamtiu et al., 2005). Different studies have also shown that using ASTs in tool support and visualization can help program comprehension, such as when learning how to code (Agrahari and Chimalakonda, 2020; Egan and McDonald, 2014). In this paper, for Java we use dataset recommendation by LeClair and McMillan (2019). We use 52M to train jam (Su et al., 2023) and further extract 25k to train CodeLlaMA (Roziere et al., 2023) for practical reasons (refer to Section 3.4). We use the same testset (i.e., the 8k testset) proposed by LeClair and McMillan (2019). For C/C++, we extract 9k methods for test and 10k methods for training from a dataset proposed by Haque et al. (2021). We used this dataset because it includes 18,775 C/C++ projects and they followed LeClair and McMillan (2019) for preprocessing. For both C/C++ and Java, we used the srcML tool srcml (2025) to generate AST in XML format in keeping with recommendations by Bansal et al. (2023).
@@
Call Graph Generation Call graph is a graph that represents the relationship between each methods (Ryder, 1979). Call graphs is an important information for helping developers comprehend source code (Alanazi et al., 2021; Walunj et al., 2019; Bansal et al., 2024; Alanazi, 2021)Mabakane (2019) also showed that call graph helps to optimize the performance of the code. In this paper, we use Doxygen (doxygen, 2023) to generate call graphs for both C/C++ and Java. Doxygen is a document generation tool that can also generate call graph. For C/C++ dataset, we extract 25k methods as our testset and 20k methods as our training set from (Haque et al., 2021). We generated the call graph within a file for C/C++ to reduce the tokens of the prompt. For Java, we used the testset proposed by LeClair and McMillan (2019) and further extracted 20k methods from the rest of dataset for finetuning. We generated the callgraph from the entire project because Java does not usually have the callgraph within the file. To reduce the tokens of the prompt, we only input the caller methods in the prompt instead of the complete files.
@@
Data flow Graph Generation Data flow graph represents how one variable passes to another method or variable. Data flow graph is important because it helps programmers to do code comprehension (Sihler, 2024; Kargén and Shahmehri, 2012) and increase productivity of software developers (Al-Saiyd, 2017). In this paper, we use Joern (joern, 2025) to generate the data flow of the method. We use Joern because Joern is highly mature for Java and C/C++. For Java dataset, we used the testset recommended by LeClair and McMillan (2019) to avoid data contamination. We further extracted 15k methods other than testset as our training set. For C/C++, we extracted 8k methods as our testset and 8k methods as our training set from (Haque et al., 2021). We use the same procedure to generate data flow for both datasets.
@@ 3.4 Code Models
JAM We used Jam (Su et al., 2023) with 1,024 context length throughout out entire experiments. jam is a language model based on GPT2 architecture for various code intelligence tasks on Java. jam was pretrained with 52m Java methods with clear train/validation/test set, public available dataset, and deduplication toolkit to avoid data contamination. While jam only has 350m parameters, Su and McMillan (2024) showed in their study that it is small enough to be useful for code intelligence tasks. The advantage of this model is that this model can be run on a single consumer hardware, e.g. 16G VRAM GPU and can be used for the experiments with more controls before scaling up to larger language models.
CodeLlaMA Instruct CodeLlaMA instruct (Roziere et al., 2023) is an open-souce model released by Meta. We used CodeLlaMA instruct because it can better understand the prompt. We used the 13B model in our experiments because it can better fit into our 24G GPU (Weyssow et al., 2025) when using Qlora (Dettmers et al., 2024). We used the Qlora released in (Poludin, 2025). CodeLlaMA is the trade-off between closed-sourced models and Jam because we have access to the models although we still do not have the control on the dataset used for pretraining.
@@
GPT-4o mini GPT-4o mini is the commercial models released by OpenAI (Achiam et al., 2023). GPT-4o mini is a closed-source model namely we do not have access to the pretraining data and we send our private source code to the model through API. Because of its closed-source, we cannot avoid data contamination issues (Balloccu et al., 2024; Jiang et al., 2024b) and data privacy issues (Wu et al., 2024). Though several studies have shown the success of GPT models (Jain and Purandare, 2025; Wang et al., 2024b).
Gemini Gemini is another commercial model released by Google (Team et al., 2023, 2024) and is enhanced for multimodal capability. Gemini is also a closed-source model. We used Gemini as another closed-souce model because Gemini has been compared with OpenAI’s ChatGPT in several studies in both software engineering (Sobo et al., 2025; Qi et al., 2023; Siam et al., 2024; Su et al., 2024) and other research topics (Qianyi, 2024; Strzalkowski et al., 2024) and has shown superior than GPT in some studies (Lee et al., 2023).
@@
RQ1 jam - code sum. code sum.
- code trans. code trans.
- code gen. code gen.
codellama - code sum. code sum.
- code trans. code trans.
- code gen. code gen.
RQ2 gpt-4o - - ast gen.
- - call graph gen.
- - dataflow gen.
gemini - - ast gen.
- - call graph gen.
- - dataflow gen.
RQ3 jam ast gen. - ast gen.
@@
Table 3: Fine-tuning hyperparameters in our experiments.
jam codellama
epoch 3 3
learning rate 3.00E-05 1.00E-04
batch size 4 2
gradient accum. steps 32 16
lora_r - 64
lora_alpha - 16
lora_dropout - 0
quant_type - nf4
bits - 4
@@ 3.5 Research Methods – RQ1
Our research method for answering RQ1 is to train each open-source model to do each code development task, then evaluate each model’s performance on that task. Tables 1 and 2 shows these configurations. For example, the first row of Table 2 shows that we train jam to do code summarization, and evaluate it on that same task. The first row of Table 1 show that the dataset we use for code summarization in Java is 170k dataset in Su and McMillan (2024). However, Table 1 also shows that we did not use jam for C/C++ datasets because jam is pretrained for Java only (see Section 3.4). Our guiding principle in all fine-tuning and evaluation is to use the datasets, hyperparameters, and other experimental settings used by papers that introduced those models and datasets; our goal is to understand model capabilities rather than optimizing performance details. We used the base jam and CodeLLaMa models released by their respective papers where those papers pretrained using large datasets of code (see Section 3.4). Then we further trained those models with the training sets which were released with each dataset (see Sections 3.2 and 3.3). We use the hyperparameters recommended in by the papers accompanying each model, which we show in Table 7. Recall from Section 3.4 that we used a LoRA process for CodeLlaMA due to its size, but a full finetune process for jam.
We evaluated each configuration using the metrics and test datasets in the papers for each dataset. For code summarization, we used the metrics METEOR and USE as implemented by Su and McMillan (2024) and recommended by Haque et al. (2022). For code translation and code generation, we used Bleu score (Papineni et al., 2002) as in the dataset that we used for experiments (Iyer et al., 2018; Lu et al., 2021).
@@ 3.6 Research Methods – RQ2
Our research method for answering RQ2 is to evaluate closed-source LLMs on static analysis tasks without any finetuning. We compare the results with the tools that we describe in Section 3.3. We show this configuration in Table 2 For example, we show that we evaluate three different static analysis tasks in the first row of RQ2 in Table 2. Our goal here is the same as RQ1. Therefore, we used the default parameters settings throughout the entire dataset although different parameters might have slightly different results.
For C/C++, we used the testset that we extracted to for evaluation. For Java, we used the teset proposed by LeClair and McMillan (2019) for evaluation. For srcML, we used Levenshtein based on (Decker et al., 2020). We used Jaccard similarity as our evaluation metrics for callgraph and data flow graph analysis (Gharibi et al., 2018; Blokhin et al., 2013; Kargén and Shahmehri, 2016). We also computed the pair accuracy and chain accuracy for both tasks.

### v2

@@ 3.3 Static Analysis Tasks
In this section, we discuss three static analysis tasks in our experiments, i.e. AST, dataflow graph, and callgraph generation with two different programming languages, i.e. C/C++ and Java.
@@
Abstract Syntax Tree Generation An Abstract Syntax Tree (AST) is a tree structure that captures the syntactic information of the source code. In addition to being a key intermediate representation in code compilation, ASTs are used to localize software bugs (neamtiu2005understanding), comprehend source code (sun2023abstract; welty1997augmenting), and in software evolution (neamtiu2005understanding). Different studies have also shown that ASTs help program comprehension, such as when learning how to code (agrahari2020astar; egan2014program). In this paper, for Java we use dataset recommendation by leclair2019recommendations. We use 52M to train jam (su2023language) and further extract 25k to train CodeLlaMA (roziere2023code) for practical reasons (refer to Section 3.4). We use the same testset (i.e., the 8k testset) proposed by leclair2019recommendations. For C/C++, we extract 9k methods for test and 10k methods for training from a dataset proposed by haque2021action. We used this dataset because they followed leclair2019recommendations for preprocessing. For both C/C++ and Java, we used the srcML tool (srcml) to generate AST in XML format in keeping with recommendations by bansal2023function.
@@
Callgraph Generation Callgraph is a graph that represents the relationship between each method (ryder1979constructing). Callgraphs are an important information for code comprehension source code (alanazi2021facilitating; walunj2019graphevo; bansal2024programmer; alanazi2021software). In this paper, we used Doxygen (doxygen) to generate callgraphs for both C/C++ and Java. For C/C++ dataset, we extracted 25k methods as our testset and 20k methods as our training set from (haque2021action). We generated the callgraph within a file for C/C++ to reduce the tokens of the prompt. For Java, we used the testset proposed by leclair2019recommendations and further extracted 20k methods from the rest of dataset for finetuning. We generated the callgraph from the entire project because Java does not usually have the callgraph within the file. To reduce the tokens of the prompt, we only provided caller methods in the prompt instead of the complete files.
@@
Dataflow Graph Generation Dataflow graph represents how one variable passes to another method or variable. Dataflow graph is important because it helps programmers to do code comprehension (sihler2024improving; kargén2012inputtracer) and increase productivity of software developers (alsaiyd2017source). In this paper, we used Joern (joern) to generate the data flow of the method. We used Joern because Joern is highly mature for Java and C/C++. For Java dataset, we used the testset recommended by leclair2019recommendations to avoid data contamination. We further extracted 15k methods other than testset as our training set. For C/C++, we extracted 8k methods as our testset and 8k methods as our training set from haque2021action. We used the same procedure to generate dataflow for both datasets.
@@ 3.4 Code Models
JAM We used Jam (su2023language) with 1,024 context length throughout out entire experiments. jam is a language model based on GPT2 architecture for various code intelligence tasks on Java. jam was pretrained with 52m Java methods with clear train/validation/test set, public available dataset, and deduplication toolkit to avoid data contamination. While jam only has 350m parameters, su2024distilled showed in their study that it is small enough to be useful for code intelligence tasks. The advantage of this model is that this model can be run on a single consumer hardware, e.g. 16G VRAM GPU and can be used for the experiments with more controls before scaling up to larger language models.
CodeLlaMA Instruct CodeLlaMA instruct (roziere2023code) is an open-souce model released by Meta. We used CodeLlaMA instruct because it can better understand the prompt. We used the 13B model in our experiments because it can better fit into our 24G GPU (weyssow2025exploring) when using Qlora (dettmers2024qlora). We used the Qlora released in (qlora). CodeLlaMA is the trade-off between closed-sourced models and Jam because we have access to the models although we still do not have the control on the dataset used for pretraining.
@@
GPT-4o mini GPT-4o mini is the commercial models released by OpenAI (achiam2023gpt). GPT-4o mini is a closed-source model namely we do not have access to the pretraining data and we send our private source code to the model through API. Because of its closed-source, we cannot avoid data contamination issues (balloccu2024leak; jiang2024investigating) and data privacy issues (wu2024unveiling). Though several studies have shown the success of GPT models (jain2025assessing; wang2024enhancing).
Gemini Gemini is another commercial model released by Google (team2023gemini; team2024gemini) and is enhanced for multimodal capability. Gemini is also a closed-source model. We used Gemini as another closed-souce model because Gemini has been compared with OpenAI’s ChatGPT in several studies in both software engineering (sobo2025evaluating; qi2023gemini; siam2024programming; su2024context) and other research topics (strzalkowski2024evaluation) and has shown superior than GPT in some studies (lee2023gemini).
@@
RQ2 jam ast gen. - ast gen.
@@
RQ4 codellama - code sum. ast gen.
- code sum. callgraph gen.
- code sum. dataflow gen.
- code trans. ast gen.
- code trans. callgraph gen.
- code trans. dataflow gen.
- code gen. ast gen.
- code gen. callgraph gen.
- code gen. dataflow gen.
3.5 Evaluation Metrics
In this section, we explain the evaluation metrics for three static analysis tasks and three code intelligence tasks.
Levenshtein distance Levenshtein distance calculates the similarity between two strings by computing the number of edits between two strings. In this paper, we used rapidfuzz implemented by rapidfuzz and converted the value to zero to one scale, where zero means two strings are dissimilar and one means two strings are the same.
Pair Accuracy Pair accuracy calculates the percentage of accurate edges in a predicted callgraph. For example, in a predicted callgraph with edges E=(A,B),(B,C)E={(A,B),(B,C)} and the correct edge E=(A,B)E={(A,B)}, the pair accuracy would be $0.5$ because there are two edges in the predicted callgraph and only one of the predicted edges is correct.
Chain Accuracy Chain accuracy calculates the percentage of accurate callgraphs in all predicted callgraphs. For example, in a predicted callgraph with edges E=(A,B),(B,C)E={(A,B),(B,C)} and the true callgraph with edges E=(A,B)E={(A,B)}, the chain accuracy would be 0 because there are two edges in the predicted callgraph, but there is only one edge in the true callgraph.
METEOR METEOR (banerjee2005meteor) is a metric to evaluate summarization tasks that considers the word similarity rather than exact word match only. The scale of METEOR is between 0 and 1, where zero means no similarity between two sentences and one means two sentences are the same.
USE USE (haque2022semantic) encodes the reference summary and the predicted summary to USE score and computes the cosine similarity between reference summary and the predicted summary. haque2022semantic shows that USE align better with human perspective. The scale of USE score is between -1 and 1, where -1 means there is no similarity between two summaries and 1 means two summaries are most similar.
CodeBERTScore CodeBERTScore (zhou2023codebertscore) is based on the BERTScore. CodeBERTScore encodes both natural language description and source code into a fixed length vector and computes the cosine similarity between encoded representation of the reference and the predicted tokens. In this paper, we used the package provided by zhou2023codebertscore to compute the CoderBERTScore F1.
Pass@k pass@k (chen2021evaluating) evaluates how many generated samples passes test cases, where k is the number of generated samples. In this paper, we used $k=1$ as our evaluation.
Compilation rate Unlike pass@k, compilation rate calculates how many generated samples are successfully compiled without the need to pass the test cases. The scale of the compilation rate is between 0 and 1, where zero means no generated code can be compiled and one means all of the generated code are compiled successfully.
3.6 Prompt Templates
In this section, we explain the prompt template for static analysis tasks. We used the same prompts for both Gemini and GPT models. Please note that the main purpose of this paper is to show that current LLMs are lack of reasoning skills instead of finding the prompts with the highest performance.
Template for dataflow analysis without examples We used the following prompt for dataflow analysis without examples, where {source_code} is the function for dataflow analysis. {main_ method_name} is the source of the dataflow and {sink} is the sink of the dataflow.
Given the code {source_code}, please do the dataflow analysis from source {main_method_name} to sink {sink} and print the statements, but without any explanation and markdown. Also, treat each statement as a statment instead of a block. Return the result with the template result:<statement1> -> <statement2> result:<statement1> -> <statement2>. Remember to use template as well.
Template for dataflow analysis with example We used the following prompt for dataflow analysis with the example, where {source_code} is the function for dataflow analysis. {main_ method_name} is the source of the dataflow and {sink} is the sink of the dataflow. {example_code} is the example function and {example_data_flow} is the dataflow of the example code.
example_code = ’’’
int main(int argc, char *argv[]) {
if (argc > 1 && strcmp(argv[1], "42") == 0) {
[…]
exit(42);
[…]
exit(0);
example_data_flow = ’’’
result: int main(int argc, char *argv[]) { ->
if (argc > 1 && strcmp(argv[1], "42") == 0) {
Given the example code {example_code} and its dataflow analysis result {example_data_flow} from source main to sink strcmp. Given the code {source_code}, please do the dataflow analysis from source {main_method_name} to sink {sink} and print the statements, but without any explanation and markdown. Please follow the example to generate the dataflow analysis result.
Template for AST generation without example We used the following prompt for AST generation without examples, where {source_code} is the function for AST generation.
Given the code {source_code}, please generate the complete srcml of the given code without any explanation and markdown. Please remember to follow srcml rules.
Template for AST generation with example We used the following prompt for AST generation with examples, where {source_code} is the function for AST generation. {example_code} is the example function and {example} is the srcml of the example code.
example = ’’’
<?xml version="1.0" encoding="UTF-8" standalone="yes"?><unit xmlns="http://www.srcML.org/srcML/src" revision="1.0.0" language="C" filename="test.c"><function><type><name>int</name></type> <name>main</name><parameter_list>()</parameter_list> <block>{<block_content><expr_stmt><expr><call><name>printf</name><argument_list> (<argument><expr><literal type="string">"Hello World!"</literal> </expr></argument>)</argument_list></call></expr>;</expr_stmt> <return>return <expr><literal type="number">0</literal></expr>;</return></block_content>}</block></function></unit>
example_code = ’’’
int main() {
printf("Hello World!");
return 0;
Heres the example code {example_code} and the srcml of that code {example}. Given the code {code}, please generate the complete srcml of that code. Please remember to follow the example. Heres the srcml:
Template for callgraph generation without example We used the following prompt for callgraph generation without examples, where {source_co-de} is the function for callgraph generation.
Given the example code {example_code} and example callgraph {example_callgraph} for the function calculate. Given the code {source_code}, please follow the example to generate all paths of the callgraph for method {function_name} in the format path: a->b->c etc, where -> means call. please just generate the graph without any explanation, markdown, additional information and remember the format is path: a->b->c
Template for callgraph generation with example We used the following prompt for callgraph generation with examples, where {source_code} is the function for callgraph generation. {example_code} is the source code example for generating callgraph and the {example_callgraph} is the callgraph for the {example_code}.
example_code = ’’’
int add(int a, int b) {
return a + b;
int multiply(int a, int b) {
return a * b;
int calculate(int x, int y) {
int sum = add(x, y);
int product = multiply(x, y);
return sum + product;
int main() {
int result = calculate(3, 4);
return result;
example_callgraph = ’’’
path: calculate -> add
path: calculate -> multiply
Given the example code {example_code} and example callgraph {example_callgraph} for the function calculate. Given the code {source_code}, please follow the example to generate all paths of the callgraph for method {function_name} in the format path: a->b->c etc, where -> means call. please just generate the graph without any explanation, markdown, additional information and remember the format is path: a->b->c
Template for AST generation after finetuning with code summarization We used the following prompt for AST generation after finetuning with code summarization, where {code} is the function for AST generation. {example_code} is the source code example for generating AST and the {example} is the AST for the {example_code}.
example = ’’’
<?xml version="1.0" encoding="UTF-8" standalone="yes"?><unit xmlns="http://www.srcML.org/srcML/src" revision="1.0.0" language="C" filename="test.c"><function><type><name>int</name> </type> <name>main</name><parameter_list>()</parameter_list><block>{<block_content><expr_stmt><expr><call><name>printf</name><argument_list>(<argument><expr><literal type="string">"Hello World!"</literal></expr></argument>)</argument_list></call></expr>;</expr_stmt><return>return <expr><literal type="number">0</literal></expr>;</return</block_content>}</block></function></unit>
example_code = ’’’
int main() {
printf("Hello World!");
return 0;
[INST] You have know how to describe the code. Now, heres the example code {example_code} and the srcml of that code {example}. can you follow the example to generate the srcml of the {code} without explanation -- just srcml [/INST]
Template for AST generation after finetuning with code generation We used the following prompt for AST generation after finetuning with code generation, where {code} is the function for AST generation. {example_code} is the source code example for generating AST and the {examp- le} is the AST for the {example_code}.
example = ’’’
<?xml version="1.0" encoding="UTF-8" standalone="yes"?><unit xmlns="http://www.srcML.org/srcML/src" revision="1.0.0" language="C" filename="test.c"><function><type><name>int</name></type> <name>main</name><parameter_list>()</parameter_list> <block>{<block_content><expr_stmt><expr><call><name>printf</name><argument_list>(<argument><expr><literal type="string">"Hello World!"</literal></expr></argument>)</argument_list></call></expr>;</expr_stmt><return>return <expr><literal type="number">0</literal></expr>;</return></block_content>}</block></function></unit>
example_code = ’’’
int main() {
printf("Hello World!");
return 0;
[…]
Template for AST generation after finetuning with code translation We used the following prompt for AST generation after finetuning with code translation, where {code} is the function for AST generation. {example_c- ode} is the source code example for generating AST and the {example} is the AST for the {example_code}.
example = ’’’
<?xml version="1.0" encoding="UTF-8" standalone="yes"?><unit xmlns="http://www.srcML.org/srcML/src" revision="1.0.0" language="C" filename="test.c"><function><type><name>int</name></type> <name>main</name><parameter_list>()</parameter_list> <block>{<block_content><expr_stmt><expr><call><name>printf</name><argument_list>(<argument><expr><literal type="string">"Hello World!"</literal></expr></argument>)</argument_list></call></expr>;</expr_stmt><return>return <expr><literal type="number">0</literal></expr>;</return></block_content>}</block></function></unit>
example_code = ’’’
int main() {
printf("Hello World!");
return 0;
[…]
Template for callgraph generation after finetuning with code summarization We used the following prompt for callgraph generation after finetuning with code summarization, where {code} is the function for callgraph generation and {method_name} is the name of the function.
[…]
Template for callgraph generation after finetuning with code generation We used the following prompt for callgraph generation after finetuning with code generation, where {code} is the function for callgraph generation and {method_name} is the name of the function.
[…]
Template for callgraph generation after finetuning with code translation We used the following prompt for callgraph generation after finetuning with code translation, where {code} is the function for callgraph generation. {method_name} is the name of the function.
[…]
Template for dataflow generation after finetuning with code summarization We used the following prompt for dataflow generation after finetuning with code summarization, where {example_data_flow} is the dataflow of the example code, {example_code} is the code for the example dataflow, {code} is the function for dataflow generation, {main_method_name} is the source of the dataflow, and {sink} is the sink of the dataflow.
example_code = ’’’
int main(int argc, char *argv[]) {
if (argc > 1 && strcmp(argv[1], "42") == 0) {
[…]
exit(42);
[…]
exit(0);
example_data_flow = ’’’
result: int main(int argc, char *argv[]) { -> if (argc > 1 && strcmp(argv[1], "42") == 0) { ’’’
[…]
Template for dataflow generation after finetuning with code generation We used the following prompt for dataflow generation after finetuning with code generation, where {example_data_flow} is the dataflow of the example code, {example_code} is the code for the example dataflow, {code} is the function for dataflow generation, {main_method_name} is the source of the dataflow, and {sink} is the sink of the dataflow.
example_code = ’’’
int main(int argc, char *argv[]) {
if (argc > 1 && strcmp(argv[1], "42") == 0) {
[…]
exit(42);
[…]
exit(0);
example_data_flow = ’’’
result: int main(int argc, char *argv[]) { -> if (argc > 1 && strcmp(argv[1], "42") == 0) { ’’’
[…]
Template for dataflow generation after finetuning with code translation We used the following prompt for dataflow generation after finetuning with code translation, where {example_data_flow} is the dataflow of the example code, {example_code} is the code for the example dataflow, {code} is the function for dataflow generation, {main_method_name} is the source of the dataflow, and {sink} is the sink of the dataflow.
example_code = ’’’
int main(int argc, char *argv[]) {
if (argc > 1 && strcmp(argv[1], "42") == 0) {
[…]
exit(42);
[…]
exit(0);
example_data_flow = ’’’
result: int main(int argc, char *argv[]) { -> if (argc > 1 && strcmp(argv[1], "42") == 0) {
[…]
@@ 3.7 Research Methods – RQ1
Our research method for answering RQ1 is to evaluate closed-source LLMs on static analysis tasks without any finetuning. We compare the results with the tools that we describe in Section 3.3. Our goal is to understand model capabilities rather than optimizing performance details. To this end, we used the default parameter settings and simple prompts throughout the entire dataset although different parameters might have slightly different results. As an additional experiment, we followed kojima2022large to add “Let’s think step by step“ at the end of our in-context learning prompt. kojima2022large observed that we can simply add “Let’s think step by step“ at the end of the prompt to improve the models’ performance on reasoning tasks.
For C/C++, we used the testset that we extracted to for evaluation. For Java, we used the testset proposed by leclair2019recommendations for evaluation. For srcML, we used Levenshtein based on (decker2020srcdiff). We used Jaccard similarity as our evaluation metrics for callgraph and data flow graph analysis (gharibi2018code2graph; blokhin2013malware; kargén2016towards). We also computed the pair accuracy and chain accuracy for both tasks.
@@ 3.8 Research Methods – RQ2
Our research method for answering RQ2 is to finetune the open-source LLMs and evaluate those models on static analysis tasks. We show those configurations in Tables 1 and 2. For example, we show that we finetune jam to do ast generation and evaluate jam with the same task. We do not finetune jam for callgraph generation because of the context length limit (see Section 3.4). Note that we increase the number of tokens to 10k for callgraph generation because callgraph generation would require larger context size. We use the hyperparameters recommended by the papers accompanying each model, which we show in Table 3. Recall from Section 3.4 that we used a QLoRA process for CodeLlaMA due to its size, but a full finetune process for jam.